# Combining Evidence Across Filtrations

### v4

@@ Abstract
In sequential anytime-valid inference, any admissible procedure must be based on e-processes: generalizations of test martingales that quantify the accumulated evidence against a composite null hypothesis at any stopping time. This paper proposes a method for combining e-processes constructed in different filtrations but for the same null. Although e-processes in the same filtration can be combined effortlessly (by averaging), e-processes in different filtrations cannot because their validity in a coarser filtration does not translate to a finer filtration. This issue arises in sequential tests of randomness and independence, as well as in the evaluation of sequential forecasters. We establish that a class of functions called adjusters can lift arbitrary e-processes across filtrations. The result yields a generally applicable “adjust-then-combine” procedure, which we demonstrate on the problem of testing randomness in real-world financial data. Furthermore, we prove a characterization theorem for adjusters that formalizes a sense in which using adjusters is necessary. There are two major implications. First, if we have a powerful e-process in a coarsened filtration, then we readily have a powerful e-process in the original filtration. Second, when we coarsen the filtration to construct an e-process, there is a logarithmic cost to recovering validity in the original filtration. [^1] Code to reproduce all empirical results is publicly available at https://github.com/yjchoe/CombiningEvidenceAcrossFiltrations.
@@ 1 Introduction
Given a null hypothesis over a sequence of observations, how do we combine evidence processes that are constructed in different filtrations? This is a recurring question in problems such as randomness testing (Vovk, 2021b, ; Ramdas et al.,, 2022), independence testing (Podkopaev et al.,, 2023; Henzi and Law,, 2024), and forecast evaluation (Henzi and Ziegel,, 2022; Choe and Ramdas,, 2023), for which the evidence processes may be constructed in coarsened filtrations. This paper addresses the problem of combining arbitrary evidence processes in such scenarios, specifically when the evidence is measured as an e-process (Ramdas et al.,, 2020).
E-processes (defined soon below) are the sequential generalization of e-values (Wasserman et al.,, 2020; Vovk and Wang,, 2021; Grünwald et al.,, 2024) and betting scores (Shafer,, 2021). An e-value $\tilde{E}$ for a composite null hypothesis $\mathcal{P}$ is a nonnegative random variable whose expectation is at most one under any null distribution, that is, $\mathbb{E}_{P}[\tilde{E}]\leq 1$ for any $P\in\mathcal{P}$. E-values generalize likelihood ratios to composite and nonparametric hypothesis tests, and $\tilde{P}=\min\{1/\tilde{E},1\}$ is a p-value for $\mathcal{P}$ by Markov’s inequality. A key benefit of using e-values, as opposed to using p-values, is that we can seamlessly combine e-values across tests by averaging them, even when the e-values are arbitrarily dependent. E-processes, which generalize the validity of an e-value to arbitrary stopping times (w.r.t. a filtration), can also enjoy the same benefit whenever we combine e-processes that have the same underlying filtration.
@@ 1.1 Background: E-processes and sequential anytime-valid inference (SAVI)
Next, we say that a sequence of random variables $(X_{t})_{t\geq 0}$ is a process in a filtration $\mathbb{F}$ (or an $\mathbb{F}$-process) if it is $\mathbb{F}$-adapted, that is, $X_{t}$ is $\mathcal{F}_{t}$-measurable for each $t$. Given a distribution $P$, we say that a nonnegative $\mathbb{F}$-process $(M_{t})_{t\geq 0}$ is a test (super)martingale (Shafer et al.,, 2011) for $P$ in $\mathbb{F}$ if $M_{0}=1$ and $\mathbb{E}_{P}[M_{t}\mid\mathcal{F}_{t-1}]=M_{t-1}$ (or $\mathbb{E}_{P}[M_{t}\mid\mathcal{F}_{t-1}]\leq M_{t-1}$ for test supermartingales) for all $t\geq 1$. Given a family of distributions $\mathcal{P}$—understood as a composite null, as opposed to a point null $P$, a nonnegative $\mathbb{F}$-process is a test (super)martingale for $\mathcal{P}$ if it is a test (super)martingale for each $P\in\mathcal{P}$. A test supermartingale possesses the optional stopping property: at any $\mathbb{F}$-stopping time $\tau$, it satisfies $\mathbb{E}_{P}[M_{\tau}]\leq 1$ for all $P\in\mathcal{P}$.
An e-process (Ramdas et al.,, 2020, 2022) is a nonnegative adapted process that satisfies the optional stopping property under the null. More formally, given a family of distributions $\mathcal{P}$, we say that a nonnegative $\mathbb{F}$-process is an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{F}$ if $\mathfrak{e}_{0}\leq 1$ and,
@@
In this sense, e-processes strictly generalize test supermartingales, and there exist composite nulls $\mathcal{P}$ for which nontrivial e-processes exist while all test supermartingales (in $\mathbb{F}$) are trivial. When $\mathcal{P}$ is understood as a (composite) null hypothesis, an e-process for $\mathcal{P}$ quantifies evidence against $\mathcal{P}$, as it is not expected to be large at arbitrary stopping times unless $\mathcal{P}$ does not adequately describe the data. Following Ramdas et al., (2022), we say that an e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ is powerful (or consistent) against a family of alternatives $\mathcal{Q}$, or $\mathcal{Q}$-powerful, if for every $Q\in\mathcal{Q}\setminus\mathcal{P}$, we have $\limsup_{t\to\infty}\mathfrak{e}_{t}=\infty$, $Q$-almost surely.
Note that the definition of an e-process explicitly depends on the filtration $\mathbb{F}$ through the choice of stopping times $\tau$. We refer to the notion of validity at arbitrary ($\mathbb{F}$-)stopping times, such as (1), as ($\mathbb{F}$-)anytime-validity (Ramdas et al.,, 2020) or ($\mathbb{F}$-)safety (Grünwald et al.,, 2024). It is known that any anytime-valid testing procedure must go through an e-process, suggesting e-processes are the central objects for SAVI (Ramdas et al.,, 2020).
We provide a glossary of the key definitions and notations in Section A. For a more comprehensive overview, we refer the reader to section 7 of Ramdas and Wang, (2025).
@@ 1.2 A motivating example from financial time series modeling
We shall now introduce a practical example motivating the problem of combining e-processes constructed in different filtrations. Below, we say that a sequence of random variables $X_{1},X_{2},\dotsc$ is random (Vovk, 2021b, ) if they are independent and identically distributed (IID).
@@
A priori, the manager does not know exactly how the data may deviate from randomness. Thus, they will want to leverage (at least) two different testing procedures that achieve power under different alternatives. One such procedure is based on the universal inference (UI; Wasserman et al.,, 2020; Ramdas et al.,, 2022) e-process, denoted as $\mathfrak{e}^{\mathsf{UI}}=(\mathfrak{e}_{t}^{\mathsf{UI}})_{t\geq 0}$, which generalizes likelihood ratios to composite nulls and alternatives. Formally, the UI e-process can be written as $\mathfrak{e}_{t}^{\mathsf{UI}}=\frac{\int_{\theta\in[0,1]^{2}}Q_{\theta}(X_{1},\dotsc,X_{t})p(\theta)d\theta}{\sup_{\mu\in[0,1]}P_{\mu}(X_{1},\dotsc,X_{t})}$, where $P_{\mu}$ is the IID $\mathsf{Ber}(\mu)$ distribution; $Q_{\theta}$ is a first-order Markov distribution with transition parameters $\theta=(p_{0\to 1},p_{1\to 1})\in[0,1]^{2}$; and $p(\theta)$ is a “prior” distribution over $\theta$. Another is based on the conformal test martingale (CTM) (Vovk, 2021b, ), denoted as $\mathfrak{e}^{\mathsf{conf}}=(\mathfrak{e}_{t}^{\mathsf{conf}})_{t\geq 0}$, where $\mathfrak{e}_{t}^{\mathsf{conf}}=\prod_{i=1}^{t}[1+\lambda(s_{i}-1/2)]$ for $\lambda\in\mathbb{R}$ and “conformal p-values” $s_{1},\dotsc,s_{t}\in[0,1]$. Intuitively, these conformal p-values capture pointwise deviations from the running mean of the data. Under the null, they are IID $\mathsf{Unif}[0,1]$ with mean $1/2$; under a changepoint alternative, their mean deviates from $1/2$ (to 0 or 1). Further details about these e-processes are not central to our discussion, so we defer them to Section B.
@@
For each fixed $t$, $\mathbb{E}_{P}[\bar{m}_{t}]\leq 1$ under any $P\in\mathcal{H}_{0}^{\mathsf{IID}}$, so $\bar{m}$ is a sequence of e-values. However, is $\bar{m}$ an e-process? To answer this question, we need to clarify which filtration we are working with. The most natural candidate is the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. However, while the UI e-process $\mathfrak{e}^{\mathsf{UI}}$ is an e-process in $\mathbb{F}$, the conformal test martingale $\mathfrak{e}^{\mathsf{conf}}$ is not. In fact, $\mathfrak{e}^{\mathsf{conf}}$ is only an e-process in a sub-filtration $\mathbb{G}=(\mathcal{G}_{t})_{t\geq 0}\subsetneq\mathbb{F}$ constructed by the conformal p-values—that is, $\mathcal{G}_{t}=\sigma(s_{1},\dotsc,s_{t})$. Note that $\mathbb{G}$ is a strictly coarser filtration than $\mathbb{F}$—for example, neither the sequence of events $\{\sum_{i=1}^{t}X_{i}\geq c\}_{t\geq 0}$ nor the raw data sequence $(X_{t})_{t\geq 0}$ is adapted to $\mathbb{G}$. Thus, $\mathfrak{e}^{\mathsf{conf}}$ is only valid at arbitrary $\mathbb{G}$-stopping times but not at arbitrary $\mathbb{F}$-stopping times. This means that $\bar{m}$ is a sequence of e-values that is not an e-process in $\mathbb{F}$; even in $\mathbb{G}$, $\bar{m}$ is not an e-process because its first term, $\mathfrak{e}_{t}^{\mathsf{UI}}$, is not $\mathcal{G}_{t}$-measurable.
@@
At this point, we note that the broader takeaway remains the same even if we substituted the CTM with its improved variants in the literature, such as the “simple jumper” CTM (e.g., Vovk et al., 2022a, ) that we describe in Section B.3. Other variants of the CTM are still invalid under $\mathbb{F}$-stopping times, and it is known that CTMs that are powerful against changepoint alternatives do not achieve power against Markovian alternatives (fig. 9.8 of Vovk et al., 2022a, ). An analogous remark could be made about the UI e-process (e.g., Saha and Ramdas,, 2024).
Some recent works have noted the challenge of combining e-processes (or test martingales) across filtrations. For randomness testing, Vovk et al., 2022a (, section 9.3 in 2nd ed.) note that even the average of the same CTMs across different runs need not be a test martingale, as each one is constructed using differently randomized conformal p-values that yield different filtrations. Besides randomness testing, there are at least two other practical examples in the literature, one in sequential forecast comparison (Henzi and Ziegel,, 2022; Choe and Ramdas,, 2023) and another in independence testing (Podkopaev et al.,, 2023; Henzi and Law,, 2024); we discuss these examples in Section 5. Yet another example is testing a scale-invariant Gaussian mean, for which different e-processes exist in different filtrations (Pérez-Ortiz et al.,, 2024; Wang and Ramdas, 2025a, ). We review this example in Section C.
Motivated by these examples, we study how to combine arbitrary e-processes across different filtrations (for the same null), such that their combination is an e-process in a finer filtration. Central to our approach is the use of adjusters (Shafer et al.,, 2011; Dawid et al., 2011a, ; Dawid et al., 2011b, ; Koolen and Vovk,, 2014; Ramdas et al.,, 2022), which are closely related to p-to-e calibrators (Vovk,, 1993; Sellke et al.,, 2001; Ramdas et al.,, 2020; Vovk and Wang,, 2021), that allow us to lift an e-process from a coarser filtration into a finer one. This is somewhat surprising because these tools were originally devised for different purposes, namely to insure against the loss of evidence over time (for adjusters) and to convert p-values to e-values (for calibrators).
Even though we focused on practical motivation thus far, our main results have standalone implications in enhancing our theoretical understanding of both adjusters and e-processes. First, we establish a novel characterization of adjusters that suggests adjusters are not only sufficient but also, in a formal sense, necessary for lifting arbitrary e-processes. Second, our $\mathfrak{e}$-lifting” result implies that, unlike in the case of test martingales, there always exists a powerful e-process for a testing problem in the data filtration whenever there exists one in a coarser filtration.

### v5

@@ Abstract
In sequential anytime-valid inference, any admissible procedure must be based on e-processes: generalizations of test martingales that quantify the accumulated evidence against a composite null hypothesis at any stopping time. This paper proposes a method for combining e-processes constructed in different filtrations but for the same null. Although e-processes in the same filtration can be combined effortlessly (by averaging), e-processes in different filtrations cannot because their validity in a coarser filtration does not translate to a finer filtration. This issue arises in sequential tests of randomness and independence, as well as in the evaluation of sequential forecasters. We establish that a class of functions called adjusters can lift arbitrary e-processes across filtrations. The result yields a generally applicable “adjust-then-combine” procedure, which we demonstrate on the problem of testing randomness in real-world financial data. Furthermore, we prove a characterization theorem for adjusters that formalizes a sense in which using adjusters is necessary. There are two major implications. First, if we have a powerful e-process in a coarsened filtration, then we readily have a powerful e-process in the original filtration. Second, when we coarsen the filtration to construct an e-process, there is a logarithmic cost to recovering validity in the original filtration.
footnotetext: Accepted for publication in the Journal of the Royal Statistical Society: Series B (Statistical Methodology).footnotetext: Code to reproduce all empirical results is publicly available at: https://github.com/yjchoe/CombiningEvidenceAcrossFiltrations.
@@ 1 Introduction
Given a null hypothesis over a sequence of observations, how do we combine evidence processes that are constructed in different filtrations? This is a recurring question in problems such as randomness testing (Vovk, 2021b; Ramdas et al., 2022), independence testing (Podkopaev et al., 2023; Henzi and Law, 2024), and forecast evaluation (Henzi and Ziegel, 2022; Choe and Ramdas, 2023), for which the evidence processes may be constructed in coarsened filtrations. This paper addresses the problem of combining arbitrary evidence processes in such scenarios, specifically when the evidence is measured as an e-process (Ramdas et al., 2020).
E-processes (defined soon below) are the sequential generalization of e-values (Wasserman et al., 2020; Vovk and Wang, 2021; Grünwald et al., 2024) and betting scores (Shafer, 2021). An e-value $\tilde{E}$ for a composite null hypothesis $\mathcal{P}$ is a nonnegative random variable whose expectation is at most one under any null distribution, that is, $\mathbb{E}_{P}[\tilde{E}]\leq 1$ under any $P\in\mathcal{P}$. E-values generalize likelihood ratios to composite and nonparametric hypothesis tests, and $\tilde{P}=\min\{1/\tilde{E},1\}$ is a p-value for $\mathcal{P}$ by Markov’s inequality. A key benefit of using e-values, as opposed to using p-values, is that we can seamlessly combine e-values across tests by averaging them, even when the e-values are arbitrarily dependent. We can also easily combine e-processes when they are defined in the same underlying filtration.
@@ 1.1 Background: E-processes and sequential anytime-valid inference (SAVI)
Next, we say that a sequence of random variables $(X_{t})_{t\geq 0}$ is a process in a filtration $\mathbb{F}$ (or an $\mathbb{F}$-process) if it is $\mathbb{F}$-adapted, that is, $X_{t}$ is $\mathcal{F}_{t}$-measurable for each $t$. Given a distribution $P$, we say that a nonnegative $\mathbb{F}$-process $(M_{t})_{t\geq 0}$ is a test (super)martingale (Shafer et al., 2011) for $P$ in $\mathbb{F}$ if $M_{0}=1$ and $\mathbb{E}_{P}[M_{t}\mid\mathcal{F}_{t-1}]=M_{t-1}$ (or $\mathbb{E}_{P}[M_{t}\mid\mathcal{F}_{t-1}]\leq M_{t-1}$ for test supermartingales) for all $t\geq 1$. Given a family of distributions $\mathcal{P}$—understood as a composite null, as opposed to a point null $P$, a nonnegative $\mathbb{F}$-process is a test (super)martingale for $\mathcal{P}$ if it is a test (super)martingale for each $P\in\mathcal{P}$. A test supermartingale possesses the optional stopping property: at any $\mathbb{F}$-stopping time $\tau$, it satisfies $\mathbb{E}_{P}[M_{\tau}]\leq 1$ for all $P\in\mathcal{P}$.
An e-process (Ramdas et al., 2020, 2022) is a nonnegative adapted process that satisfies the optional stopping property under the null. More formally, given a family of distributions $\mathcal{P}$, we say that a nonnegative $\mathbb{F}$-process is an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{F}$ if $\mathfrak{e}_{0}\leq 1$ and,
@@
In this sense, e-processes strictly generalize test supermartingales, and there exist composite nulls $\mathcal{P}$ for which nontrivial e-processes exist while all test supermartingales (in $\mathbb{F}$) are trivial. When $\mathcal{P}$ is understood as a (composite) null hypothesis, an e-process for $\mathcal{P}$ quantifies evidence against $\mathcal{P}$, as it is not expected to be large at arbitrary stopping times unless $\mathcal{P}$ does not adequately describe the data. Following Ramdas et al. (2022), we say that an e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for a null hypothesis $\mathcal{P}$ is powerful (or consistent) against a family of alternatives $\mathcal{Q}$, or $\mathcal{Q}$-powerful, if for every $Q\in\mathcal{Q}\setminus\mathcal{P}$, we have $\limsup_{t\to\infty}\mathfrak{e}_{t}=\infty$, $Q$-almost surely.
Note that the definition of an e-process explicitly depends on the filtration $\mathbb{F}$ through the choice of stopping times $\tau$. We refer to the notion of validity at arbitrary ($\mathbb{F}$-)stopping times, such as (1), as ($\mathbb{F}$-)anytime-validity (Ramdas et al., 2020) or ($\mathbb{F}$-)safety (Grünwald et al., 2024). It is known that any anytime-valid testing procedure must go through an e-process, suggesting e-processes are the central objects for SAVI (Ramdas et al., 2020).
We provide a glossary of the key definitions and notations in Section A. For a more comprehensive overview, we refer the reader to section 7 of Ramdas and Wang (2025).
@@ 1.2 A motivating example from financial time series modeling
We shall now introduce a practical example motivating the problem of combining e-processes constructed in different filtrations. Below, we say that a sequence of random variables $X_{1},X_{2},\dotsc$ is random (Vovk, 2021b) if they are independent and identically distributed (IID).
@@
A priori, the manager does not know exactly how the data may deviate from randomness. Thus, they will want to leverage (at least) two different testing procedures that achieve power under different alternatives. One such procedure is based on the universal inference (UI) e-process (Wasserman et al., 2020; Ramdas et al., 2022), denoted as $\mathfrak{e}^{\mathsf{UI}}=(\mathfrak{e}_{t}^{\mathsf{UI}})_{t\geq 0}$, which generalizes likelihood ratios to composite nulls and alternatives. Formally, the UI e-process can be written as $\mathfrak{e}_{t}^{\mathsf{UI}}=\frac{\int_{\theta\in[0,1]^{2}}Q_{\theta}(X_{1},\dotsc,X_{t})p(\theta)d\theta}{\sup_{\mu\in[0,1]}P_{\mu}(X_{1},\dotsc,X_{t})}$, where $P_{\mu}$ is the IID $\mathsf{Ber}(\mu)$ distribution, $Q_{\theta}$ is a first-order Markov distribution with transition parameters $\theta=(p_{0\to 1},p_{1\to 1})\in[0,1]^{2}$, and $p(\theta)$ is a “prior” distribution over $\theta$. Another is based on the conformal test martingale (CTM) (Vovk, 2021b), denoted as $\mathfrak{e}^{\mathsf{conf}}=(\mathfrak{e}_{t}^{\mathsf{conf}})_{t\geq 0}$, where $\mathfrak{e}_{t}^{\mathsf{conf}}=\prod_{i=1}^{t}[1+\lambda(s_{i}-1/2)]$ for $\lambda\in\mathbb{R}$ and “conformal p-values” $s_{1},\dotsc,s_{t}\in[0,1]$. Intuitively, these conformal p-values capture pointwise deviations from the running mean of the data. Under the null, they are IID $\mathsf{Unif}[0,1]$ with mean $1/2$; under a changepoint alternative, their mean deviates from $1/2$ (to 0 or 1). Further details about these e-processes are not central to our discussion, so we defer them to Section B.
@@
For each fixed $t$, we see that $\mathbb{E}_{P}[\bar{m}_{t}]\leq 1$ under any $P\in\mathcal{H}_{0}^{\mathsf{IID}}$, so $\bar{m}$ is a sequence of e-values. However, is $\bar{m}$ an e-process? To answer this question, we first need to clarify which filtration we are working with. The most natural candidate is the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(X_{1},\dotsc,X_{t})$. While $\mathfrak{e}^{\mathsf{UI}}$ is an e-process in $\mathbb{F}$, $\mathfrak{e}^{\mathsf{conf}}$ is not. In fact, $\mathfrak{e}^{\mathsf{conf}}$ is only an e-process in a sub-filtration $\mathbb{G}=(\mathcal{G}_{t})_{t\geq 0}$ of $\mathbb{F}$ constructed by the conformal p-values—that is, $\mathcal{G}_{t}=\sigma(s_{1},\dotsc,s_{t})$. This sub-filtration $\mathbb{G}$ is strictly coarser than $\mathbb{F}$ (e.g., neither the event sequence $\{\sum_{i=1}^{t}X_{i}\geq c\}_{t\geq 0}$ nor the raw data sequence $(X_{t})_{t\geq 0}$ is adapted to $\mathbb{G}$). Thus, $\mathfrak{e}^{\mathsf{conf}}$ is only valid at arbitrary $\mathbb{G}$-stopping times but not at arbitrary $\mathbb{F}$-stopping times (i.e., not $\mathbb{F}$-anytime-valid), implying that $\bar{m}$ is a sequence of e-values that is not an e-process in $\mathbb{F}$. Even in the coarse filtration $\mathbb{G}$, $\bar{m}$ is not an e-process because it is not adapted to $\mathbb{G}$.
@@
At this point, we note that the broader takeaway remains the same even if we substituted the CTM with its improved variants in the literature, such as the “simple jumper” CTM (e.g., Vovk et al., 2005) that we describe in Section B.3. Other variants of the CTM are still invalid under $\mathbb{F}$-stopping times, and it is known that CTMs that are powerful against changepoint alternatives do not achieve power against Markovian alternatives (fig. 9.8 of Vovk et al., 2005). An analogous remark could be made about the UI e-process (e.g., Saha and Ramdas, 2024).
Some recent works have noted the challenge of combining e-processes (or test martingales) across filtrations. For randomness testing, Vovk et al. (2005, section 9.3 in 2nd ed.) note that even the average of the same CTMs across different runs need not be a test martingale, as each one is constructed using differently randomized conformal p-values that yield different filtrations. Besides randomness testing, there are at least two other practical examples in the literature, one in sequential forecast comparison (Henzi and Ziegel, 2022; Choe and Ramdas, 2023) and another in independence testing (Podkopaev et al., 2023; Henzi and Law, 2024); we discuss these examples in Section 5. Yet another example is testing a scale-invariant Gaussian mean, for which different e-processes exist in different filtrations (Pérez-Ortiz et al., 2024; Wang and Ramdas, 2025a). We review this example in Section C.
Motivated by these examples, we study how to combine arbitrary e-processes across different filtrations (for the same null), such that their combination is an e-process in a finer filtration. Central to our approach is the use of adjusters (Shafer et al., 2011; Dawid et al., 2011a, b; Koolen and Vovk, 2014; Ramdas et al., 2022), which are closely related to p-to-e calibrators (Vovk, 1993; Sellke et al., 2001; Ramdas et al., 2020; Vovk and Wang, 2021), that allow us to lift an e-process from a coarser filtration into a finer one. This is somewhat surprising because these tools were originally devised for different purposes, namely to insure against the loss of evidence over time (for adjusters) and to convert p-values to e-values (for calibrators).
Although we focused on practical motivation thus far, our main results have standalone implications in enhancing our theoretical understanding of both adjusters and e-processes. First, we establish a novel characterization of adjusters that suggests adjusters are not only sufficient but also, in a formal sense, necessary for lifting arbitrary e-processes. Second, our $\mathfrak{e}$-lifting” result implies that, unlike in the case of test martingales, there always exists a powerful e-process for a testing problem in the data filtration whenever there exists one in a coarser filtration.